OpenStudy (anonymous):

Can someone explain Pythagorean Theorem to me a bit more clearly? It's kind of confusing.

pooja195 (pooja195):

This should help :) https://www.mathsisfun.com/pythagoras.html

pooja195 (pooja195):

\(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Huge\bf\color{#FF0000}W\color{#FF4900}e\color{#FF9200}l\color{#FFDB00}c\color{#FFff00}o\color{#B6ff00}m\color{#6Dff00}e\color{#24ff00}~\color{#00ff00}t\color{#00ff49}o\color{#00ff92}~\color{#00ffDB}O\color{#00ffff}p\color{#00DBff}e\color{#0092ff}n\color{#0049ff}S\color{#0000ff}t\color{#2400ff}u\color{#6D00ff}d\color{#B600ff}y\color{#FF00ff}!\)\(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Large\bf\color{Lime}{Please,\ don't\ hesitate\ to\ ask\ questions!}\)\(\Large\bf\color{magenta}{We're\ here\ to\ help!}\Large\color{magenta}{\ddot\smile}\)

OpenStudy (anonymous):

OH MA GHERD

OpenStudy (anonymous):

WTAF

OpenStudy (anonymous):

DA RAINBOW DOOOOEEE

OpenStudy (anonymous):

Basically, with any right triangle (a triangle with a 90 degree angle), the length of one leg, a, squared, plus the length of the other leg, b, squared, equals the length of the big leg (hypotenuse), c, squared.

jimthompson5910 (jim_thompson5910):

The pythagorean theorem is \(\Large a^2 + b^2 = c^2\) 'a' and 'b' are the sides of the legs of the right triangle c is the hypotenuse. It's always the longest side and always opposite the 90 degree angle The term \(\Large a^2\) means "a squared" The term \(\Large b^2\) means "b squared" The term \(\Large c^2\) means "c squared" When you square a number, you are multiplying it by itself. Visually in terms of geometry, you are finding the area of a square with that given side length Notice how in this image http://mathandmultimedia.com/wp-content/uploads/2010/02/pythagorean2.png they have attached squares to the yellow triangle. It turns out that if you add the areas of the green and red squares, you get the blue square's area So `green square + red square = blue square` more or less This is another way of stating \(\Large a^2 + b^2 = c^2\) This video shows that in action https://www.youtube.com/watch?v=CAkMUdeB06o Note: each square in the video is really a prism, but each prism has the same height so it's analogous to what I'm saying above